OpenStudy (loser66):

\[\int_{-3}^{3} |x+1|dx\] Please, help

OpenStudy (jtvatsim):

Integrate separately on parts of the domain that are convenient is the quick tip. :)

OpenStudy (loser66):

Show me, please.

OpenStudy (jtvatsim):

Or, just draw it and find the area. :)

OpenStudy (jtvatsim):

Will do, I'll do the drawing method since it's more clever.

OpenStudy (loser66):

I show you my confusion.

OpenStudy (jtvatsim):



OpenStudy (loser66):

no confuse, hehehe.. I am sorry. I posted the wrong question. But I would like to know how to solve it traditionally

OpenStudy (jtvatsim):

No worries. :) You should be able to see from the graph that the answer is the area of two triangles. But I will do the traditional (algebraic nastiness) approach as well.

OpenStudy (loser66):

I meant integrate separately..........

OpenStudy (jtvatsim):

First, we need the definition of what |x + 1| means: |x + 1| = { x + 1, for x >= -1; -(x + 1), for x < -1. This is the usual, "make the number positive" rule we remember in our heads expressed algebraically.

OpenStudy (jtvatsim):

It may take a bit of pondering to fully see that. :)

OpenStudy (jtvatsim):

Then, this naturally gives us two convenient domains. (-infinity, -1) and (-1, +infinity). The split point is at x = -1.

OpenStudy (jtvatsim):

We want (-3, 3) so we break this into the domains (-3, -1), and (-1, 3).

OpenStudy (jtvatsim):

Then we integrate using the definition of absolute value.

OpenStudy (jtvatsim):

\[\int\limits_{-3}^3 |x + 1| \ dx = \int\limits_{-3}^{-1} -(x+1) \ dx + \int\limits_{-1}^3 x+1 \ dx\]

OpenStudy (jtvatsim):

The integration can then be carried out as usual.

OpenStudy (loser66):

Got you. Thanks a lot

OpenStudy (jtvatsim):

No problems! :)